OpenStudy (anonymous):

Leroy has $15 less in his pocket than Donna has. If Donna had $8 more, she would have 4 times as much money as Leroy.

OpenStudy (lgbasallote):

let x = Leroy's money y = Donna's money x = y-15 y + 8 = 4(y-15) y + 8 = 4y - 60 3y = 68 y = $22.66 x = $7.66

OpenStudy (anonymous):

x=y-15 y-8=4x